OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4. 2 3 All real numbers No solution

OpenStudy (anonymous):

i don't get this

OpenStudy (sophiaholder):

OK so what is the first things you think you should do?

OpenStudy (sophiaholder):

thing*

OpenStudy (sophiaholder):

Try thinking of them as two seperate equations you need to simplify.

OpenStudy (anonymous):

All real numbers is the solution To simplify: −2x−6=−2x−6 Then you have to add 2x to both sides:−2x−6+2x=−2x−6+2x −6=−6 Then add: −6+6=−6+6 0=0